If the lengths of a pair of opposite sides of a rectangle were doubled and the other pair of opposite sides were each decreased by 14 m, the new rectangle would be a square with a perimeter of 84 m. Find the rectangle's original dimensions. ........
Answer by privani(34)

Understood from question:
Square's perimeter = 84m, with sides "s"
Hence,
4s = 84
s = 84/4
s = 21
Further from question:
"a pair of opposite sides of a rectangle were doubled"; assuming this side is "x"
2x = 21
x = 21/2
x = 10.5
"the other pair of opposite sides were each decreased by 14 m"; assuming the other side is "y"
y - 14 = 21
y = 21 + 14
y = 35
So, the rectangle's original dimension is 10.5 m x 35 m.